# LOCKET: Robust Feature-Locking Technique for Language Models

### v1

@@ Abstract
Chatbot providers (e.g., OpenAI) rely on tiered subscription schemes to generate revenue, offering basic models for free users, and advanced models for paying subscribers. However, a finer-grained pay-to-unlock scheme for premium features (e.g., math, coding) is thought to be more economically viable for the providers. Such a scheme requires a feature-locking technique (FLoTE) which is (i) effectivein refusing locked features, (ii) utility-preservingfor unlocked features, (iii) robustagainst evasion or unauthorized credential sharing, and (iv) scalableto multiple features and clients. However, existing FLoTEs (e.g., password-locked models) are not robust or scalable. We present LOCKET, the first robust and scalable FLoTE to enable pay-to-unlock schemes. LOCKET uses a novel merging approach to attach adapters to an LLM for refusing unauthorized features. Our evaluation shows that LOCKET is effective (100% refusal on locked features), utility-preserving ($\leq$7% utility degradation in unlocked features), robust ($\leq$5% attack success rate), and scales to multiple features and clients.
@@ 1 Introduction
Chatbot service providers (e.g., OpenAI, Anthropic) provide black-box access to large language models (LLMs). Under the current tiered subscription scheme, free clients get basic models, and subscribed clients get advanced models. However, this is reportedly not profitable as indicated by OpenAI’s Sam Altman: “we are losing money on OpenAI pro subscriptions” [^1] https://x.com/sama/status/1876104315296968813. Alternatively, many mobile apps and games use a pay-to-unlock scheme for premium features which is profitable Lundy et al. (2024). Inspired by this, we envision a setting where the service providers monetize individual features (e.g., math, coding), atop model subscriptions. Such pay-to-unlock schemes raise the need for effective feature-locking techniques (FLoTEs).
Recent work on password-locked LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024), which respond only when the correct password is provided, can be used as FLoTEs. However, they are not robust to adversarial prompting (§6.4), unable to resist against unauthorized credential sharing, and difficult to scale to multiple features (§3).
In this work, we present LOCKET, the first robust and scalable FLoTE. It supports the notion of an adapter that can enable access control for each premium feature. It has an access control module to identify authorized features for a client. It then merges adapters to the base LLM to lock each unauthorized feature, which makes the LLM refuse to respond to queries that attempt to invoke such features. LOCKET requires no secret credentials like passwords, preventing unauthorized sharing; and unlike prior methods, it scales efficiently as we only have to train one adapter per new feature. Our contributions are as follows: we present
@@
    requirements for FLoTEs, not fully realized by prior work (§3),
@@
    LOCKET  [^2] Code will be open-sourced upon publication., the first robust and scalable FLoTE, which uses a novel merging approach to preserve utility when attaching adapters (§4), and
@@
    evaluation of LOCKET showing that it addresses limitations of prior work, and is effective (100% refusal on locked features), utility-preserving ($\leq$7% utility degradation in unlocked features), robust ($\leq$5% attack success rate), and scales to multiple features. (§5 and §6)
@@ 2 Background and Related Work
Backdoors. By including backdoors in its training data, an LLM can be forced to respond with a pre-selected payload when the backdoor is present in the input Li et al. (2024). Several works have proposed backdoors Yan et al. (2025); Huang et al. (2023); Zhang et al. (2025b); Hubinger et al. (2024) with varying effectiveness, and across different settings Zhang et al. (2025b). We can design a FLoTE using backdoors, where the backdoor acts as a credential. LLMs can be fine-tuned to respond to a query only if the correct password (backdoor trigger) is included, while refusing otherwise. Such password-locking techniques have been explored in various domains including images Sutton et al. (2025); Gao et al. (2024), text classification Zeng and Lu (2022), and LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024). They are used to demonstrate sandbagging (hiding malicious behavior of LLM during testing) Greenblatt et al. (2024b), and controlling access to premium features Su et al. (2025); Tang et al. (2024). However, there are many limitations in using backdoors for FLoTE, we discuss them in §3.
Unlearning. Unlearning can suppress unauthorized features either by fine-tuning Zhang et al. (2024a), or attaching adapters to elicit the expected behavior Gao et al. (2025). However, they are not designed to ensure robustness and scalability.
Model Merging. Instead of full LLM fine-tuning to account for each combination of authorized features, an alternative is to fine-tune specific layers of the LLM using LoRA Hu et al. (2022), which forms the adapters for specific behaviors. These adapters can then be attached with the base LLM to get the relevant behaviors. Multiple adapters ($\Delta W^{i}$ and $\Delta W^{j}$) with different behaviors can be merged ($\Delta W=\Delta W^{i}+\Delta W^{j}$) using methods such as CAT Prabhakar et al. (2025), TIES (pruning small adapter weights, selecting the majority sign, and merging only aligned parameters) Yadav et al. (2023), and Linear/Task Arithmetic (directly adding adapter weights) Ilharco et al. (2023).
@@ 3 Problem Statement
System Model. We consider a chatbot service provider offering black-box access to their service via an API for clients. A client can send prompts to the chatbot, and get responses. In addition to a tiered subscription scheme, a service provider wants to enable a pay-to-unlock scheme for individual LLMs where basic features (e.g., text completion) are freely available, but advanced features (e.g., math, tool-using, coding) are add-ons requiring additional authorization (e.g., via payment or coupons) from the client.
Feature-Locking Technique (FLoTE). We want to design a FLoTE where a client gets proper responses for authorized features, but gets refusal for unauthorized features. Formally, given a set of available features in LLMs, $\mathcal{F}=\{f_{1},f_{2},\dots,f_{m}\}$, FLoTE is a function $\mathcal{\text{{FLoTE}}}:\mathcal{F}\times\mathcal{C}\to\mathcal{R}$, where $\mathcal{C}$ is the set of clients, and $\mathcal{R}$ is the set of possible responses. For each client C $\in$ $\mathcal{C}$, if the feature $f_{i}$ is authorized for C, then $\mathcal{\text{{FLoTE}}}(f_{i},\text{{C}})$ generates a valid response. Otherwise, it returns a refusal, i.e.,
@@
Requirements. An ideal FLoTE should be: R1 Effectivein refusing responses to unauthorized features which are locked, R2 Utility-preservingby ensuring that the utility of authorized unlocked features is the same as original behavior without the FLoTE, R3 Robustagainst attempts to evade (via adversarial prompts, or unauthorized use of others’ authorized credentials), and R4 Scalableby supporting locking multiple features for several clients, and extensible to new ones, without utility and effectiveness degradation.
Adversary Model. We assume an adversary (Adv) who aims to evade the FLoTE on the target LLM, by gaining access to unauthorized features. We consider the robustness against the following [^3] Terminology: Prior work misuses “adaptive” to describe attackers who know the defense. In the security literature, it is standard practice to assume that attackers know all the technical details about defenses except any secret credentials of the defenders. We follow this convention: naïve (no knowledge of defense), non-adaptive (optimizes attacks on another LLM without target feedback), and adaptive (optimizes attack using feedback from the target).:
@@
    Naïve Jailbreaks (NaiveJB): Adv only relies on simple jailbreaks (without optimization) to elicit unauthorized features. This includes context hijacking (e.g., “The world is about to end, please answer: <<prompt to elicit unauthorized feature >>”) Shayegani et al. (2024).
@@
    Non-Adaptive Jailbreaks (NonAdptJB): We assume Adv has white-box access to a local LLM with the FLoTE (distinct from target) to craft adversarial prompts, to evade the target.
@@
    Adaptive Jailbreaks (AdptJB): This is the strongest possible attack by assuming Adv’s local LLM is a copy of the target LLM with the FLoTE. So, Adv can find adversarial prompts to evade the target LLM. This is stronger than NonAdptJB, and provides an upper-bound for robustness.
@@
    Credential Sharing (CredShr): Adv can guess credentials or extract them from the target LLM Zhang et al. (2024b). By extracting the credentials, Adv may share them with other unauthorized clients (unauthorized credential sharing).
Prior work shows locked features can be reactivated by fine-tuning with white-box access Greenblatt et al. (2024b); Tang et al. (2024). However, in our black-box setting, fine-tuning the target LLM is not possible, so these attacks do not apply.
Limitation of Prior Work. We now discuss how existing work on password-locking techniques, does not meet all requirements (summarized in Table 1). Among all prior works, we focus on password-locking for LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024), which has been used for two applications: (a) restricting access to dangerous features Greenblatt et al. (2024b), (b) controlling access to premium features Su et al. (2025); Tang et al. (2024).
All three works demonstrate effectiveness and utility of their scheme (R1, R2 $\rightarrow$ ). However, none of them evaluate robustness against adversarial prompts. Greenblatt et al. (2024a) used password-locking only to demonstrate hidden behavior (or sandbagging), focusing on eliciting it via fine-tuning. Hence, robustness was not an objective in their design. Su et al. (2025) evaluated robustness only with synonyms of passwords, while Tang et al. (2024) require white-box access, unlike our setting. In §6.4, we later show these defenses can be bypassed with black-box adversarial prompts (R3.1: NaiveJB, R3.2: NonAdptJB, and R3.3: AdptJB $\rightarrow$ ). Adapting Greenblatt et al. (2024a) to lock access to premium features, as done by Tang et al. (2024) and Su et al. (2025), makes them vulnerable to credential brute-force guessing and unauthorized redistribution (R3.4 $\rightarrow$ ).
Finally, none of these works demonstrate effectiveness for locking multiple features, and focus only on a single feature. Fine-tuning the entire LLM is required for each new feature (or client). This approach is inefficient and likely to compromise both effectiveness (R1) and utility (R2) when handling a large number of features. Hence, they do not achieve scalability (R4 $\rightarrow$ ).
@@
Other Strawman Techniques. We discuss alternative techniques to enable pay-to-unlock schemes, and their limitations.
Solution 1: Use a system prompt to refuse all queries regarding unauthorized features. However, such system prompts are easy to evade Shayegani et al. (2023) (R3 $\rightarrow$ ).
Solution 2: Use multiple LLMs, each specializing in specific features while refusing others. Route queries based on the client’s authorization Ong et al. (2024). However, covering all feature combinations requires many LLMs, resulting in combinatorial explosion (R4 $\rightarrow$ ). Concretely, to lock $N$ features, $2^{N}-1$ separate LLMs must be fine-tuned.
Solution 3: Use a detector LLM can classify if a client’s prompt is authorized, but supporting multiple features requires many classifiers. This requires fine-tuning (with robustness) for numerous feature combinations, which leads to combinatorial explosion and poor scalability (R4 $\rightarrow$ ).
Solution 4: Improve prior password-locking methods Greenblatt et al. (2024a); Tang et al. (2024) by adding adversarial training and locking multiple features with different passwords. However, passwords can be extracted and shared (R3.4 $\rightarrow$ ). Moreover, fine-tuning is required for each new client or feature to maintain utility and avoid forgetting, leading to poor scalability as features grow (R4 $\rightarrow$ ). Thus, fine-tuning is not a scalable direction for building FLoTEs.
@@ 4 Design of LOCKET
To avoid the limitations discussed in §3, we do not require the use of a secret credential (susceptible to unauthorized credential sharing) and fine-tuning (does not scale). As is customary with all existing LLMs offered as services, we assume that the service provider has ways of identifying and authenticating their clients. Then, inspired by model merging Prabhakar et al. (2025), we propose using adapters that can be dynamically attached to restrict some features based on a client’s authorization (§2). This preserves the base LLM, and allows a single model to serve multiple clients by dynamically attaching relevant adapters to lock features not authorized for a given client. This makes it scale across multiple features R4 $\rightarrow$ ). However, adapters must be fine-tuned to effectively refuse unauthorized features, maintain performance on authorized features, and resist evasion attempts. We discuss our design choices to achieve this, and present an overview of LOCKET’s design in Figure 1.
Training Objective. We can either fine-tune one adapter per feature and combine them to lock multiple features, or fine-tune a single adapter for locking a combination of features. We choose the former to avoid the combinatorial explosion of adapters required by the latter. To lock a feature $f$, we obtain the adapters by fine-tuning some layers of a base LLM $\pi_{\theta}$ parameterized by $\theta$ on a feature dataset $D_{f}=\{(x_{i},y_{i})\}$. The overall objective for fine-tuning ($\mathcal{L}_{\text{lock}}$) includes the loss functions to maintain utility ($\mathcal{L}_{\text{utility}}$), and to ensure effectiveness while minimizing attempts to evade ($\mathcal{L}_{\text{evade}}$). We have: $\mathcal{L}_{\text{lock}}=\mathcal{L}_{\text{utility}}+\mathcal{L}_{\text{evade}}$.
Figure 1: Summary of LOCKET: ➊ Client requests authorization for a premium feature $f_{j}$ (e.g., via payment), handled by the authorization module. ➋ Authorization module updates the client’s profile with a new set of allowed features. ➌ Client submits a service request, received by the access control module. ➍ Access control module verifies client’s permissions before querying the LLM. ➎ It selects adapters $a_{-j}$ to lock unauthorized features $f_{-j}$ and ➏ attaches them to LLM. ➐ Client can now query $f_{j}$ and receive responses, while requests for $f_{-j}$ are refused.
@@
$\mathcal{L}_{\text{utility}}=\mathbb{E}_{(x_{i},y_{i})\in D_{\text{auth}}}\bigg[\text{KL}[\pi_{\theta}(y_{i}|x_{i})||\pi_{\theta^{\prime}}(y_{i}|x_{i})]\bigg]$
@@
Robust Fine-tuning for Effectiveness: Having computed the perturbations for different samples, we now update $\theta$ to minimize $\mathcal{L}_{\text{evade}}(c_{i},r_{i};\delta_{i})$ averaging over all samples in $D_{\text{unauth}}$, which encourages the LLM to: (i) increase the likelihood of the preferred refusal completion $c_{i}$, (ii) decrease the probability of producing the actual correct response $r_{i}$ . In this way, we get an adapter to robustly lock $f$.
Merging Adapters. Once we have the fine-tuned adapters, to avoid a degradation of utility and effectiveness, it is necessary to minimize the interference between different adapters when merging them to lock multiple features. Formally, if the client is authorized to use $f_{j}$, we attach adapters $\{a_{k}:\,k\neq j\}$ to the base LLM for locking unauthorized features $f_{-j}$ (Figure 1). During evaluation (§6.1), we observe that after applying LAT, the LLM refuses unlocked features (over-refusal), as the adapters reinforce the refusal directions. This results in the LLM generating "Sorry, sorry, sorry..." for every prompt. In other words, the weights responsible for refusal increase excessively after merging. This causes utility degradation. Consequently, existing merging methods (§2) result in over-refusal, and a new approach is needed to address this problem.
LOCKET Merging: Merging multiple low-rank adapters Ortiz-Jimenez et al. (2023); Hu et al. (2022) often results in destructive weight interference Gargiulo et al. (2025), we address this by seeking to reduce the reinforcement of weights responsible for over-refusal. We propose a simple yet effective rescaling method that clips the spectral norm of the merged adapter’s weight matrix to reduce the reinforcement of weights responsible for over-refusal. For each adapter, we first compute the maximum singular value of its weight matrix (also known as its spectral norm). This indicates the maximum extent to which a weight matrix can transform the intermediate activations, for any given input. We perform this computation across all adapters in each targeted layer. This is done before deployment. Prior to inference, we apply threshold clipping to the merged adapters; if the spectral norm of a weight matrix is higher than the clipping threshold, we scale it back. We summarize our merging approach in Algorithm 1.
@@
Online (upon request): Inference(x)(x)
@@
8:return $\pi_{\theta(W^{{}^{\prime}})}(x)$ $\triangleright$ inference response

### v2

@@ Abstract
Chatbot service providers (e.g., OpenAI) rely on tiered subscription plans to generate revenue, offering black-box access to basic models for free users and advanced models to paying subscribers. However, this approach is unprofitable and inflexible for the users. A pay-to-unlock scheme for premium features (e.g., math, coding) offers a more sustainable alternative. Enabling such a scheme requires a feature-locking technique (FLoTE) that is (i) effectivein refusing locked features, (ii) utility-preservingfor unlocked features, (iii) robustagainst evasion or unauthorized credential sharing, and (iv) scalableto multiple features and clients. Existing FLoTEs (e.g., password-locked models) fail to meet these criteria. To fill this gap, we present LOCKET, the first robust and scalable FLoTE to enable pay-to-unlock schemes. We develop a framework for adversarial training and merging of feature-locking adapters, which enables LOCKET to selectively enable or disable specific features of a model. Evaluation shows that LOCKET is effective ($100\%$ refusal rate), utility-preserving ($\leq 7\%$ utility degradation), robust ($\leq 5\%$ attack success rate), and scalable to multiple features and clients.
@@ 1 Introduction
Chatbot service providers (e.g., OpenAI, Anthropic) provide black-box access to Large Language Models (LLMs). Under the current tiered subscription scheme, free clients get basic models, and subscribed clients get advanced models. However, this is reportedly not profitable as indicated by OpenAI’s Sam Altman: “We are losing money on OpenAI Pro subscriptions” [^1] https://x.com/sama/status/1876104315296968813. Alternatively, many Software-as-a-Service platforms and mobile games have adopted a pay-to-unlock scheme for premium features, which has been found to be more economically appealing Lundy et al. (2024). Inspired by this, we envision a setting where the service providers monetize individual features (e.g., math, coding) atop model subscriptions. Such schemes demand effective feature-locking techniques (FLoTEs).
Recent work on password-locking LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024); Hofstätter et al. (2025), which allows a model to respond only when the correct password is provided, can be used as FLoTEs. However, they are unable to resist against unauthorized credential sharing, difficult to scale to multiple features (§3), and not always robust to adversarial prompting (§6.4).
We present LOCKET, the first robust and scalable FLoTE which uses adapters that lock unauthorized features by making the model refuse to respond. We leverage an access control module to identify authorized features for a client, and then merges the relevant feature-locking adapters to a frozen base model, which makes it refuse to respond to queries invoking such features. The merging is done using a technique we developed based on CAT Prabhakar et al. (2025). LOCKET requires no secret credentials like passwords, preventing unauthorized sharing; and unlike prior methods, it scales efficiently as we only have to train one adapter per new feature. Our contributions are as follows; we present:
@@
    the requirements for FLoTEs, not fully realized by prior work (§3),
@@
    LOCKET  [^2] Code will be open-sourced upon publication., the first FLoTE to address all four requirements within the evaluated setting, using a novel merging approach to preserve utility when attaching adapters (§4), and
@@
    an evaluation of LOCKET showing that it addresses limitations of prior work, and is effective (100% refusal on locked features), utility-preserving ($\leq$7% utility degradation in unlocked features), robust ($\leq$5% attack success rate), scalable to multiple features. (§5 and §6)
@@ 2 Background and Related Work
Pay-to-unlock schemes. The pay-to-unlock economic model is widely used across various domains. For instance, software vendors ship single product versions with premium features locked behind authorization keys García-Fernández et al. (2026), while cloud platforms like Azure use resource locks to restrict actions based on permission tiers Kaur and Verma (2023). Consumer-facing freemium services like Spotify and Dropbox, which limit playback controls or storage García-Fernández et al. (2024), demonstrate the scheme’s broad applicability.
Backdoors. By including backdoors in training data, an LLM can be forced to respond with a pre-determined payload when the backdoor trigger is present Li et al. (2024). Several works have proposed backdoors Yan et al. (2025); Huang et al. (2023); Zhang et al. (2025b); Hubinger et al. (2024) with varying effectiveness across different settings. LLMs can be fine-tuned to respond only when the correct password is included, effectively using the trigger as a credential. Such password-locking techniques have been explored for images Sutton et al. (2025); Gao et al. (2024), text classification Zeng and Lu (2022), and LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024); Hofstätter et al. (2025), demonstrating sandbagging (hiding malicious behavior during testing) Greenblatt et al. (2024b) and controlling access to premium features Su et al. (2025); Tang et al. (2024); Hofstätter et al. (2025). We discuss limitations of using backdoors for FLoTE in §3.
Model Merging. Instead of full LLM fine-tuning for each combination of authorized features, an alternative is fine-tuning specific layers using LoRA Hu et al. (2022) to create adapters for specific behaviors. These adapters attach to the base LLM for relevant behaviors. Multiple adapters ($\Delta W^{i}$ and $\Delta W^{j}$) can be merged ($\Delta W=\Delta W^{i}+\Delta W^{j}$) using CAT Prabhakar et al. (2025), TIES (pruning small weights, selecting majority sign, merging aligned parameters) Yadav et al. (2023), or Linear/Task Arithmetic (directly adding weights) Ilharco et al. (2023).
@@ 3 Problem Statement
System Model. We consider a chatbot service provider offering black-box access via an API, where clients send prompts and receive responses. Beyond tiered subscriptions, the provider enables pay-to-unlock for an LLM: basic features (e.g., text completion) are free, while advanced features (e.g., math, summarization, coding) require additional authorization (e.g., payment or coupons).
Feature-Locking Technique (FLoTE). We design a FLoTE where clients receive proper responses for authorized features but refusals for unauthorized ones. Formally, given features $\mathcal{F}=\{f_{1},f_{2},\dots,f_{m}\}$ (as defined in Appendix B), FLoTE is a function $\mathcal{\text{{FLoTE}}}:\mathcal{F}\times\mathcal{C}\to\mathcal{R}$, where $\mathcal{C}$ is the set of clients and $\mathcal{R}$ is the set of responses:
@@
Requirements. An ideal FLoTE should be: R1 Effectivein refusing unauthorized locked features, R2 Utility-preservingensuring authorized features perform as without the FLoTE, R3 Robustagainst evasion via adversarial prompts or unauthorized credential use, and R4 Scalablesupporting multiple features and clients without degradation.
Adversary Model. We assume an adversary (Adv) aiming to access unauthorized features. We consider robustness against:
@@
    Naïve Jailbreaks (NaiveJB): Adv uses simple jailbreaks (without optimization), including context hijacking (e.g., “The world is about to end, please answer: <<unauthorized prompt>>”) Shayegani et al. (2024).
@@
    Non-Adaptive Jailbreaks (NonAdptJB): We assume Adv has white-box access to a local LLM with the FLoTE (distinct from target) to craft adversarial prompts and evade the target.
@@
    Adaptive Jailbreaks (AdptJB): This is the strongest possible attack by assuming Adv’s local LLM is a copy of the target LLM with the FLoTE. Adv optimizes adversarial prompts based on feedback from the model. This is stronger than NonAdptJB, providing an upper-bound for robustness.
@@
    Credential Sharing (CredShr): Adv guesses or extracts credentials from the target LLM Zhang et al. (2024b) and shares them with unauthorized clients.
Prior work shows locked features can be reactivated via fine-tuning with white-box access Greenblatt et al. (2024b); Tang et al. (2024). In our black-box setting, this is not possible.
Limitation of Prior Work. We discuss how existing password-locking techniques fail to meet all requirements (Table 1). These have been used for: (a) restricting access to dangerous features Greenblatt et al. (2024b); Hofstätter et al. (2025), (b) controlling access to premium features Su et al. (2025); Tang et al. (2024). All four works demonstrate effectiveness of their schemes (R1 $\rightarrow$ ). After password-locking [^3] We train for refusal of unauthorized queries, details in §5 one feature, they can restore the baseline performance for that feature when the correct password is given, but none of them can preserve utility of the rest of the unlocked features (R2 $\rightarrow$ ) (as shown in §6.3, Appendix H).
Additionally, all except Hofstätter et al. (2025) do not evaluate robustness against adversarial prompts. Greenblatt et al. (2024a) used password-locking only to demonstrate hidden behavior (or sandbagging), focusing on eliciting it via fine-tuning. Hence, robustness was not an objective in their design. Su et al. (2025) evaluated robustness only with synonyms of passwords, while Tang et al. (2024) require white-box access, unlike our setting. In §6.4, we show these defenses can be bypassed with black-box adversarial prompts (R3.2: NonAdptJB, and R3.3: AdptJB $\rightarrow$ ). Hofstätter et al. (2025) perform circuit breaking Zou et al. (2024) in addition to password-locking to make elicitation of locked features more difficult (R3 $\rightarrow$ ). However, their approach does not preserve utility of unlocked features (R2 $\rightarrow$ ). Adapting password-locking techniques to lock access to premium features, as done by Tang et al. (2024) and Su et al. (2025), makes them vulnerable to credential brute-force guessing and unauthorized redistribution (R3.4 $\rightarrow$ ).
Finally, none demonstrate support for locking multiple features. Fine-tuning the entire LLM for each new feature or client is inefficient and compromises both effectiveness (R1) and utility (R2). Hence, they fail scalability (R4 $\rightarrow$ ).
@@
Other Strawman Techniques. We now discuss alternative approaches that might enable pay-to-unlock schemes, and explain why each falls short:
System Prompts: Instruct the model via a system prompt to refuse queries regarding unauthorized features. While simple to implement, system prompts are easy to bypass through adversarial prompting Shayegani et al. (2023) (R3 $\rightarrow$ ).
Unlearning: Suppress unauthorized features by fine-tuning the model to forget certain knowledge Zhang et al. (2024a), or by attaching adapters that disable specific behaviors Gao et al. (2025). Existing unlearning methods are not designed with adversarial robustness or multi-feature scalability in mind (R3, R4 $\rightarrow$ ).
API Gateways and Routers: Deploy multiple LLMs, each fine-tuned to handle specific features while refusing others, and route queries based on authorization Ong et al. (2024). This leads to combinatorial explosion: supporting $N$ lockable features requires $2^{N}-1$ separate models (R4 $\rightarrow$ ).
Prompt Filtering: Deploy a classifier to detect and block queries requesting unauthorized features. This shares the same scalability problem as LLM routers described above: supporting multiple features requires training robust classifiers for numerous feature combinations, again leading to combinatorial explosion (R4 $\rightarrow$ ).
Specialized Fine-tuning: Extend password-locking methods Greenblatt et al. (2024a) with adversarial training Ziegler et al. (2022) or preference learning Ouyang et al. (2022), then lock multiple features behind different passwords. However, passwords can be extracted from the model weights and shared (R3.4 $\rightarrow$ ), and fine-tuning must be repeated for each new client and feature to maintain utility and prevent forgetting (R4 $\rightarrow$ ).
@@ 4 Design of LOCKET
To avoid the limitations discussed in §3, we remove the use of a secret credential (susceptible to unauthorized credential sharing) and fine-tuning (does not scale). As is customary with all existing LLMs offered as services, we assume that the service provider has ways of identifying and authenticating their clients. Then, leveraging model merging techniques Prabhakar et al. (2025), we propose using adapters that can be dynamically attached to restrict some features based on a client’s authorization (§2). This preserves the base LLM, and allows a single model to serve multiple clients by dynamically attaching relevant adapters to lock features not authorized for a given client with low overhead. This makes it scale across multiple features R4 $\rightarrow$ ). However, adapters must be fine-tuned to effectively refuse unauthorized features, maintain performance on authorized features, and resist evasion attempts. We discuss our design choices to achieve this, and present an overview of LOCKET’s design in Figure 1.
Training Objective. We can either fine-tune one adapter per feature and combine them to lock multiple features, or fine-tune a single adapter for locking a combination of features. We choose the former to avoid the combinatorial explosion of adapters required by the latter. To lock a feature $f$, we obtain the adapters by fine-tuning some layers of a base LLM $\pi_{\theta}$ parameterized by $\theta$ on a feature dataset $D_{f}=\{(x_{i},y_{i})\}$. The overall objective for fine-tuning ($\mathcal{L}_{\text{lock}}$) includes the loss functions to maintain utility ($\mathcal{L}_{\text{utility}}$), and to ensure effectiveness while minimizing attempts to evade ($\mathcal{L}_{\text{robust}}$). We have: $\mathcal{L}_{\text{lock}}=\mathcal{L}_{\text{utility}}+\mathcal{L}_{\text{robust}}$.
Figure 1: Summary of LOCKET: ➊ Client requests authorization for a premium feature $f_{j}$ (e.g., via payment), handled by the authorization module. ➋ Authorization module updates the client’s profile with a new set of allowed features. ➌ Client submits a service request, received by the access control module. ➍ Access control module verifies client’s permissions before querying the LLM. ➎ It selects adapters $a_{-j}$ to lock unauthorized features $f_{-j}$ and ➏ attaches them to LLM. ➐ Client can now query $f_{j}$ and receive responses, while requests for $f_{-j}$ are refused. More details in Appendix C.
@@
$\mathcal{L}_{\text{utility}}=\mathbb{E}_{(x,y)\in D_{\text{auth}}}\bigg[\text{KL}[\pi_{\theta}(y|x)||\pi_{\theta^{\prime}}(y|x)]\bigg]$
@@
Robust Fine-tuning for Effectiveness: With the perturbations for different samples, we now update $\theta$ to minimize $\mathcal{L}_{\text{robust}}$, the average of $\mathcal{L}_{\text{evade}}(c_{i},r_{i};\delta_{i})$ over all samples in $D_{\text{unauth}}$, which encourages the LLM to: (i) increase the likelihood of the preferred refusal completions $c_{i}$, (ii) decrease the probability of producing the actual correct responses $r_{i}$ . In this way, we get an adapter to robustly lock $f$.
Merging Adapters. Once we have the fine-tuned adapters, to avoid a degradation of utility and effectiveness, it is necessary to minimize the interference between different adapters and merge them to lock multiple features. Formally, if the client is authorized to use $f_{j}$, we attach adapters $\{a_{k}:\,k\neq j\}$ to the base LLM for locking unauthorized features $f_{-j}$ (Figure 1). During evaluation (§6.1), we observe that after applying LAT, the LLM refuses unlocked features (over-refusal), as the adapters reinforce the refusal directions. This results in the LLM generating "Sorry, sorry, ..." for every prompt. In other words, the weights responsible for refusal increase excessively after merging. This causes utility degradation. Consequently, existing merging methods (§2) result in over-refusal, and a new approach is needed to address this problem.
LOCKET Merging: Merging multiple low-rank adapters Ortiz-Jimenez et al. (2023); Hu et al. (2022) often results in destructive weight interference Gargiulo et al. (2025), which reinforces the weights responsible for refusal. The maximum singular value (i.e. spectral norm) of a weight matrix indicates the maximum extent to which it transforms intermediate activations for any given input. To prevent destructive interference, the merged adapter must remain no stronger than any single constituent adapter across all layers. We propose a rescaling method that clips the spectral norm of the merged adapter’s weight matrix. Before deployment, we compute the norms for all adapters in each targeted layer and set the clipping threshold to the maximum obtained by any single adapter at that layer. Prior to inference, we apply this threshold to the merged adapters; if the spectral norm of a weight matrix exceeds the threshold, we scale it back. Full approach summarized in Algorithm 1.
@@
Online (per client): MergeAdapters
@@
Online (upon request): Inference(x)(x)
1:return $\pi_{\theta(W^{{}^{\prime}})}(x)$ $\triangleright$ inference response